OpenStudy (anonymous):

What 3 positive numbers give the same result when multiplied and added together?

OpenStudy (anonymous):

The simplest answer is 1 and 2 and 3. 1+2+3=6 1x2x3=6 Other valid answers: 0, 0, 0 -1, 0, 1 -x, 0, x 1, 1.5, 5 The general formula here is: abc=a+b+c abc-a=b+c a(bc-1)=b+c a=(b+c)/(bc-1) If you choose two random numbers b and c you will ALWAYS have a number a that satisfies the conditions (unless bc=1)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

You have to 25%

OpenStudy (anonymous):

well, thats what it says